# Graph Reinforcement Learning for Combinatorial Optimization: A Survey and Unifying Perspective

### v1

@@ 1 Introduction
Figure 1: Visual summary of the structure and topics of the present survey.
@@
In recent years, Machine Learning (ML) has started to emerge as a valuable tool in approaching combinatorial optimization problems, with researchers in the field anticipating its impact to be transformative (Bengio et al., 2021; Cappart et al., 2021). Most notably, the paradigm of Reinforcement Learning (RL) has shown the potential to discover, by trial-and-error, algorithms that can outperform traditional exact methods and (meta)heuristics. A common pattern is to express the problem of interest as a Markov Decision Process (MDP), in which an agent incrementally constructs a solution, and is rewarded according to its ability to optimize the objective function. Starting from the MDP formulation, a variety of RL algorithms can be transparently applied, rendering this approach very flexible in terms of the problems it can address. In parallel, works that address graph combinatorial optimization problems with RL have begun to emerge in a variety of scientific disciplines spanning chemistry (You et al., 2018a), computer science (Valadarsky et al., 2017), economics (Darvariu et al., 2021b), and statistics (Zhu et al., 2020), to name but a few.
@@
The goal of this survey is to present a unified framework, which we term Graph RL, for decision-making problems over graphs. We will synthetize a variety of approaches that can be interpreted in the context of this emerging paradigm. We will discuss several combinatorial optimization problems with a focus on non-canonical ones, for which efficient, performant algorithms are not typically known. Indeed, recent surveys have focused on works that apply RL to canonical problems, a term we use to refer to problems which have been been intensely studied, possibly for decades. For example, research on solving the aforementioned TSP alone dates back nearly 70 years to the paper of Dantzig et al. (1954), and very effective algorithms exist for solving the problem optimally (Applegate et al., 2009) or approximately (Lin & Kernighan, 1973; Helsgaun, 2000) for instances with up to tens of millions of nodes. Other notable examples of canonical problems addressed in the RL literature include Maximum Independent Sets (Ahn et al., 2020), Maximum Cut (Khalil et al., 2017; Ahn et al., 2020), as well as routing problems such as the Vehicle Routing Problem (VRP) (Kool et al., 2019; Kim & Park, 2021). With a few exceptions, even though work on such benchmark problems is important for pushing the limitations of ML-based methods, currently they cannot directly compete with well-established, highly optimized heuristic and exact solvers. This article is therefore complementary to other surveys (Mazyavkina et al., 2021; Wang & Tang, 2021) and perspectives (Bengio et al., 2021; Cappart et al., 2021) on RL for combinatorial optimization, both in terms of proposing a unifying paradigm and its focus on non-canonical problems.
@@ 2 Background
2.1 Graph Fundamentals
Graphs, also called networks, are the underlying mathematical objects that are the focus of the present survey.111 Regarding the difference between the terms, “graph” is more accurately used to refer to the mathematical abstraction, while “network” refers to a realization of this general concept, such as a particular social network. The terms are synonymous in general usage (Barabási, 2016, Chapter 2.2) and we use them interchangeably in the remainder of this work. Their origin is in graph theory, the genesis of which is considered to be Euler’s work on the bridges of Königsberg, which poses the problem of finding a route through a city that crosses each bridge exactly once (Biggs et al., 1986). The field has since developed as a branch of discrete mathematics, and solutions have been developed for problems such as route-finding (Doran & Michie, 1966), coloring (Jensen & Toft, 1995), and graph enumeration (Harary & Palmer, 1973).
The related modern discipline of network science (Newman, 2018) draws on concepts from a variety of fields in order to perform fundamental studies of graph-structured systems, builds predictive models and algorithms, and applies them to practical problems. To name but a few examples: probability theory and statistical physics have been used to analytically examine the properties and limit behavior of networks (Barrat et al., 2008); graph algorithms and data structures, with origins in theoretical computer science, find broad uses when using network modeling (Even, 2011); insights into networks can be leveraged for optimization problems that are relevant in an operational context, such as supply and logistics chains (Ahuja et al., 1995).
Let us fix some key notation. We denote a graph $G$ as the tuple $(V,E)$, where $V$ is a set of nodes or vertices that are used to represent the entities that are part of the system, and $E$ is a set of edges that represent connections and relationships between the entities. We indicate an element of the set $V$ with $v$ or $v_{i}$ and an element of $E$ with $e$ or $e_{i,j}$, with the latter indicating the edge between the nodes $v_{i}$ and $v_{j}$.
A common mathematical structure used for representing graphs in computer programs is the adjacency matrix $\mathbf{A}$, in which the entry at row $i$ and column $j$ is equal to $1$ if an edge exists between nodes $v_{i}$ and $v_{j}$, and 00 otherwise. Since many networks are sparse in terms of the number of connections per node, in practical implementations an edge list representation, which only records existing edges, may be used instead. If $\mathbf{A}$ is symmetric, which implies reciprocal bidirectional connectivity, the graph is known as undirected. Otherwise, it is called directed.
@@ 2.1 Graph Fundamentals
Nodes and edges may optionally have attribute vectors associated with them, which we denote as $\mathbf{x}_{v}$ and $\mathbf{x}_{e}$ respectively. These can capture various aspects about the problem of interest depending on the application domain, and may be either static or dynamic. Some examples include geographical coordinates in a space, the on-off status of a node, and the capacity of an edge for the transmission of information or a physical quantity. Equipped with such attributes, graphs become a powerful mathematical tool for studying a variety of systems.
2.2 Combinatorial Optimization on Graphs
@@
Methods from network science allow us to formally characterize processes taking place over a graph. For example, decision-makers might be interested in the global structural properties such as the efficiency with which the network exchanges information, or its robustness when network elements fail, aspects crucial to infrastructure networks (Latora & Marchiori, 2001; Albert et al., 2000). One can also use the graph formalism to model flows of quantities such as packets or merchandise, relevant in a variety of computer and logistics networks (Ahuja, 1993). Taking a decentralized perspective, we may be interested in the individual and society-level outcome of  network games, in which a network connects individuals that take selfish decisions in order to maximize their gain (Jackson & Zenou, 2015).
@@
  • Exact methods: approaches that solve the problem exactly, i.e., will find the globally optimal solution if it exists. These include exact search algorithms such as brute force search. Notably, if the problem of interest has a linear objective, one can formulate it as an (integer) linear program (Thie & Keough, 2011), for which efficient solving methods such as the simplex method (Dantzig & Thapa, 1997) and branch-and-bound (Land & Doig, 1960) exist. Typically, exact methods only work well for small to medium-sized problem instances.
  • Heuristics (Pearl, 1984) and approximation algorithms (Williamson & Shmoys, 2011): approaches that do not guarantee to find the optimal solution, but instead find one in a best-effort fashion. For the latter category, one can also obtain theoretical guarantees on the approximation ratio between the obtained solution and the optimal one. Such approaches make use of insights about the structure of the problem and objective function at hand, and can typically scale to large problem instances.
@@
2.3 Complexity of Combinatorial Optimization Problems
We have stated that combinatorial optimization problems are challenging, motivating the existence of heuristic and metaheuristic approaches for solving them. How might we make more precise statements about the difficulty of a problem? Computational complexity theory offers a possible answer. In this subsection, we briefly discuss some key concepts that are relevant for Graph Reinforcement Learning, given they often arise in the literature but are only mentioned in passing. A formal treatment is out of scope of the present work, and we refer interested readers to (Garey & Johnson, 1979), which gives a catalogue of problems and an extensive treatment of the area, and (Goldreich, 2008), which is a more didactic and accessible resource.
The complexity of problems is typically analyzed in terms of resource requirements in space and time that an algorithm requires to solve them in relation to input size. A complexity class is a set of problems that share similar requirements. Decision problems have a binary (yes or no) answer, and are most commonly the subject of analysis.
The $\mathcal{NP}$ complexity class consists of the decision problems for which, given a candidate solution, its correctness can be verified in polynomial time with respect to the inputs. The $\mathcal{P}$ complexity class is the subset of $\mathcal{NP}$ containing the problems that can also be solved in polynomial time. Whether problems that are efficiently verifiable are also efficiently solvable (i.e., whether $\mathcal{P}\stackrel{{\scriptstyle?}}{{=}}\mathcal{NP}$) is a longstanding unsolved problem in computer science.
A technique for relating the complexity of problems to one another is to transform them via polynomial-time reductions. To solve problem A, we can map its inputs (via a reduction that takes polynomial time) to problem B, and use an algorithm to solve B. This means that A is "no harder to solve" than B, implying a series of relationships between the difficulty of problems. Those problems that are at least as difficult to solve as all other problems in $\mathcal{NP}$ are called $\mathcal{NP}$-complete. In other words, a problem is $\mathcal{NP}$-complete if (1) candidate solutions can be verified in polynomial time and (2) all problems in $\mathcal{NP}$ can be reduced to it in polynomial time.
The pioneering work of Richard M. Karp (Karp, 1972) demonstrated the $\mathcal{NP}$-completeness of 21 decision problems by constructing polynomial-time reductions from the Boolean satisfiability problem (SAT). This relies on the Cook-Levin theorem (Cook, 1971; Levin, 1973), another cornerstone result in theoretical computer science, which establishes the $\mathcal{NP}$-completeness of SAT. The work of Karp gave a blueprint for reasoning about the difficulty of problems, yielding a reductibility hierarchy between them, with SAT at the root. Karp mentions that the theorems “suggest, but do not imply, that these problems, as well as many others, will remain intractable perpetually”.
We note that, while such tools are useful constructs to formally categorize the difficulty of decision problems, it may be impractical to solve them by conversion to a set of Boolean clauses and use of a SAT solver. The reduction procedure may be inefficient (e.g., a high-degree polynomial), and better algorithms that take advantage of the specific problem structure can be designed.
We also remind the reader that all statements in this subsection up to now have been specifically about decision problems. Let us return to the running TSP example: its $\mathcal{NP}$-completeness can be established by reduction from Undirected Hamiltonian Cycle, one of Karp’s 21 problems. The decision variant of TSP asks whether there exists a tour such that its total length is at most a certain bound $b$, which has a binary answer.
Instead, the optimization variant of TSP asks to find the tour with optimal length. In this context, it is unclear how we might satisfy condition (1) for $\mathcal{NP}$-completeness (i.e., verify a candidate solution in polynomial time), given that we would need to know the optimal tour length to begin with. Problems that satisfy condition (2) above but not (1) are known as $\mathcal{NP}$-hard. In the present survey, we are typically concerned with optimization problems, which are the more practical of the two variants. Decision problems, instead, are more amenable to this type of mathematical analysis.
@@ 2.3 Complexity of Combinatorial Optimization Problems
Let us close this subsection with a few remarks related to the emerging works in Graph Reinforcement Learning covered by this survey. Complexity can be characterized formally under specific settings and assumptions, and authors have referred to prior complexity results to justify the approximate methodology employed due to the difficulty of a problem. For example, graph construction with the objective of maximizing algebraic connectivity (Mosk-Aoyama, 2008), causal discovery with discrete random variables (Chickering et al., 2004), and determining weights for shortest path routing over graphs (Fortz & Thorup, 2004) have all been shown to be $\mathcal{NP}$-hard.
@@
More broadly, however, the formulations considered (which may be flexible with respect to the considered reward function or include learned models) can render general statements about the problem difficulty impossible. Due to this, many works proceed with formulating and solving the optimization problem heuristically without formally establishing its complexity class. Despite this, the problems are clearly difficult in a practical sense due to their combinatorial nature. Given the lack of more general tools to characterize complexity appropriately, authors often resort to describing the sizes of the state and action spaces or the time complexity of the RL agent to decide an action with respect to input size.
@@ 2.4 Machine Learning for Combinatorial Optimization
In recent years, ML has started to emerge as a valuable tool in approaching combinatorial optimization problems, with researchers in the field anticipating its impact to be transformative (Bengio et al., 2021; Cappart et al., 2021). Worthy of note are the following relationships and “points” of integration at the intersection of ML and combinatorial optimization:
@@
    ML can enable the discovery of new algorithms through the use of Reinforcement Learning (RL), another ML paradigm. Broadly speaking, RL is a mechanism for producing goal-directed behavior through trial-and-error (Sutton & Barto, 2018). In this framework, one formulates the problem of interest as a Markov Decision Process (MDP), which can be solved in a variety of ways. In the RL paradigm, an agent interacts by means of actions with an uncertain environment, receiving rewards that are proportional to the optimality of its actions; the objective of the agent is to adjust its behavior so as to maximize the sum of rewards received.
@@ 2.5.2 Dimensions of RL Algorithms
One important distinction is that between model-based algorithms (which assume knowledge of the MDP) and model-free algorithms (which require only samples of agent-environment interactions). To be specific, the state space $\mathcal{S}$ and action space $\mathcal{A}$ are assumed to be known; a model $\mathcal{M}=(P,R)$ refers to knowing, or having some estimate of, the transition and reward functions $P,R$. Model-based methods can incorporate knowledge about the world to greatly speed up learning. The model can either be given a priori or learned. In the former case, it typically takes the form of a set of mathematical descriptions that fully define $P$ and $R$. In the latter case, learning $P$ corresponds to a density estimation problem, while learning $R$ is a Supervised Learning problem. Learning architectures for this purpose can range from probabilistic models such as Gaussian Processes (Deisenroth et al., 2011) to deep neural networks (Oh et al., 2015). Model-based RL is especially advantageous where real experience is expensive to generate and executing poor policies may have a negative impact (e.g., in robotics). When equipped with a model of the world, the agent can plan its policy, either at the time actions need to be taken focusing on the current state (decision-time planning), or not focused on any particular state (background planning). Model-free algorithms, on the other hand, can yield simpler learning architectures. This comes at the expense of higher sample complexity: they require more environment interactions to train. The two categories can also be combined: it is possible to use a model to generate transitions, to which a model-free algorithm can be applied (Gu et al., 2016). It is worth noting that, in the context of graph combinatorial optimization, we can often describe the transition and reward functions analytically based on the process of interest. For the TSP, both the transition and reward functions can be expressed in closed form and do not involve any uncertainty induced by the environment or other actors. Hence, in case the underlying model $\mathcal{M}$ is known, model-based algorithms can be leveraged.

### v2

@@ 1 Introduction
Figure 1: Visual summary of the structure and topics of the present survey. $\mathcal{G}$ and $\mathcal{K}$ denote the sets of possible graph structures and graph control actions, respectively; $\mathcal{F}$ is a real-valued objective function that serves as the optimization target. The goal for Structure Optimization is to find the optimal structure $G$, while Process Optimization involves finding a set of optimal control actions $\kappa$.
@@
In recent years, Machine Learning (ML) has started to emerge as a valuable tool in approaching combinatorial optimization problems, with researchers in the field anticipating its impact to be transformative (Bengio et al., 2021; Cappart et al., 2023). Most notably, the paradigm of Reinforcement Learning (RL) has shown the potential to discover, by trial-and-error, algorithms that can outperform traditional exact methods and (meta)heuristics. A common pattern is to express the problem of interest as a Markov Decision Process (MDP), in which an agent incrementally constructs a solution, and is rewarded according to its ability to optimize the objective function. Starting from the MDP formulation, a variety of RL algorithms can be transparently applied, rendering this approach very flexible in terms of the problems it can address. In parallel, works that address graph combinatorial optimization problems with RL have begun to emerge in a variety of scientific disciplines spanning chemistry (You et al., 2018a), computer science (Valadarsky et al., 2017), economics (Darvariu et al., 2021b), and statistics (Zhu et al., 2020), to name but a few.
@@
The goal of this survey is to present a unified framework, which we term Graph RL, for combinatorial decision-making problems over graphs. Indeed, recent surveys have focused on works that apply RL to canonical problems, a term we use to refer to problems which have been intensely studied, possibly for decades. For example, research on solving the aforementioned TSP alone dates back nearly 70 years to the paper of Dantzig et al. (1954), and very effective algorithms exist for solving the problem optimally (Applegate et al., 2009) or approximately (Lin & Kernighan, 1973; Helsgaun, 2000) for instances with up to tens of millions of nodes. Other notable examples of canonical problems addressed in the RL literature include Maximum Independent Sets (Ahn et al., 2020), Maximum Cut (Khalil et al., 2017; Ahn et al., 2020), as well as routing problems such as the Vehicle Routing Problem (VRP) (Kool et al., 2019; Kim & Park, 2021). With a few exceptions, even though work on such benchmark problems is important for pushing the limitations of ML-based methods, currently they show inferior performance to well-established, highly optimized heuristic and exact solvers.
In contrast, there have been a number of works in recent years that approached non-canonical graph combinatorial optimization problems with RL techniques, showing that RL can achieve superior performance over non-RL methods in empirical evaluations. For these problems, performant exact or approximate algorithms are not known, and (meta)heuristic algorithms are typically leveraged. This is due to the fact that these are less-studied problems, but also because, in many cases, they are harder to formalize using existing techniques. [^1] These two characteristics are intertwined in a sense: indeed, if the problems are difficult to formalize, it is then more difficult to apply existing classes of techniques or use tools like solvers.To name but a few examples: Yang et al. (2023b) surpassed the performance of hill climbing, simulated annealing, greedy search, and an evolutionary algorithm for structural network rewiring (Section 3.2); Darvariu et al. (2021b) outperformed simulated annealing, best response, a payoff transfer approach, and other heuristics based on local node properties for finding equilibria in network games (Section 4.2); Meirom et al. (2021) achieved improvements compared to local node heuristics and the Local Index Rank (LIR) algorithm (Liu et al., 2017) for influence maximization (Section 4.3); Shen et al. (2018) surpassed several widely used methods that learn vector representations of entities and relations for knowledge base completion (Section 4.4). A synthesis of the high-level reasons for the better comparative performance of RL techniques on these problems is given in Section 6.1.
Therefore, in this article, we systematize the variety of approaches that comprise the emerging Graph RL paradigm. We set out to elucidate the landscape of similarities and differences in problem formulations, RL algorithms, and function approximation techniques that have been adopted, since they can differ substantially. In introducing this framework, we notice that works can naturally be divided depending on whether the goal is to optimize the structure of graphs or the outcome of processes taking place over them, dedicating Sections 3 and 4 to these topics respectively. We note that our work is complementary to other surveys (Mazyavkina et al., 2021; Wang & Tang, 2021) and perspectives (Bengio et al., 2021; Cappart et al., 2023) on RL for combinatorial optimization, both in terms of proposing a unifying paradigm and its focus on non-canonical problems, which these works have largely ignored.
Inclusion criteria. For a work to be covered by this survey, we require that (1) the problem under study is fully defined as an abstract graph combinatorial optimization problem with a discrete search space and the goal of optimizing an objective function defined over the solution space, (2) that the problem does not have existing satisfactory exact or approximate algorithms, and (3) that the problem is addressed by casting it in the MDP framework and solved approximately using an RL method, interpreted broadly to include planning and imitation learning. Therefore, examples of works that we do not treat are: those that use a graph representation as part of a larger pipeline for solving a particular application scenario (e.g., as is common in robotics and multi-agent systems); work on canonical problems such as the TSP and VRP for which effective algorithms already exist; and papers that use approaches falling outside of the MDP framework.
@@ 2 Background
2.1 Graphs and Combinatorial Optimization
@@
Graphs, also called networks, are the underlying mathematical objects that are the focus of the present survey.222 Regarding the difference between the terms, “graph” is more accurately used to refer to the mathematical abstraction, while “network” refers to a realization of this general concept, such as a particular social network. The terms are synonymous in general usage (Barabási, 2016, Chapter 2.2) and we use them interchangeably in the remainder of this work. We denote a graph $G$ as the tuple $(V,E)$, where $V$ is a set of nodes or vertices that are used to describe the entities that are part of the system, and $E$ is a set of edges that represent connections and relationships between the entities. We indicate an element of the set $V$ with $v$ or $v_{i}$ and an element of $E$ with $e$ or $e_{i,j}$, with the latter indicating the edge between the nodes $v_{i}$ and $v_{j}$. The adjacency matrix is denoted by $\mathbf{A}$. Nodes and edges may optionally have attribute vectors associated with them, which we denote as $\mathbf{x}_{v}$ and $\mathbf{x}_{e}$ respectively. These can capture various aspects of the problem of interest depending on the application domain, and may be either static or dynamic. Some examples include geographical coordinates in a space, the on-off status of a node, and the capacity of an edge for transmitting information or a physical quantity. Equipped with such attributes, graphs become a powerful mathematical tool for studying a variety of systems.
@@
Methods from network science (Newman, 2018) allow us to formally characterize processes taking place over a graph. For example, decision-makers might be interested in the global structural properties such as the efficiency with which the network exchanges information, or its robustness when network elements fail, aspects crucial to infrastructure networks (Latora & Marchiori, 2001; Albert et al., 2000). One can also use the graph formalism to model flows of quantities such as packets or merchandise, relevant in a variety of computer and logistics networks (Ahuja, 1993). Taking a decentralized perspective, we may be interested in the individual and society-level outcome of  network games, in which a network connects individuals that take selfish decisions in order to maximize their gain (Jackson & Zenou, 2015).
@@
  • Exact methods: approaches that solve the problem exactly, i.e., will find the globally optimal solution if it exists. Notably, if the problem of interest has a linear objective, one can formulate it as an (integer) linear program (Thie & Keough, 2011), for which efficient solving methods such as the simplex method (Dantzig & Thapa, 1997) and branch-and-bound (Land & Doig, 1960) exist.
  • Heuristics (Pearl, 1984) and approximation algorithms (Williamson & Shmoys, 2011): approaches that do not guarantee to find the optimal solution, but instead find one in a best-effort fashion. For the latter category, one can also obtain theoretical guarantees on the approximation ratio between the obtained solution and the optimal one. Such approaches make use of insights about the structure of the problem and objective function at hand. They can typically scale to larger problem instances than exact methods.
@@
Many combinatorial optimization problems are $\mathcal{NP}$-hard, motivating the existence of heuristic and metaheuristic approaches for solving them. A formal treatment is out of scope of the present work, and we refer interested readers to (Garey & Johnson, 1979), which gives a catalogue of problems and an extensive treatment of the area, and (Goldreich, 2008), which is a more didactic and accessible resource. For the emerging works in Graph Reinforcement Learning covered by this survey, complexity can occasionally be characterized formally under specific settings and assumptions; where possible, researchers have referred to prior complexity results to justify the approximate methodology employed due to the difficulty of a problem. For example, graph construction with the objective of maximizing algebraic connectivity (Mosk-Aoyama, 2008), causal discovery with discrete random variables (Chickering et al., 2004), and determining weights for shortest path routing over graphs (Fortz & Thorup, 2004) have all been shown to be $\mathcal{NP}$-hard.
@@
More broadly, however, the formulations considered, which may include learned models, can render general statements about the problem very difficult or impossible. Furthermore, the goal of many RL approaches is to accommodate different reward functions: the complexity of the resulting problems is highly dependent on the choice of the latter. Due to this, many works proceed with formulating and solving the optimization problem approximately without formally establishing its complexity class. Given the lack of more general tools to characterize complexity appropriately, authors often resort to describing the sizes of the state and action spaces or the time complexity required by the RL agent to decide an action with respect to input size.
@@ 2.2 Machine Learning for Combinatorial Optimization
In recent years, ML has started to emerge as a valuable tool in approaching combinatorial optimization problems, with researchers in the field anticipating its impact to be transformative (Bengio et al., 2021; Cappart et al., 2023). Worthy of note are the following relationships and “points” of integration at the intersection of ML and combinatorial optimization:
@@
    ML can enable the discovery of new algorithms through the use of Reinforcement Learning (RL), another ML paradigm. Broadly speaking, RL is a mechanism for producing goal-directed behavior through trial-and-error (Sutton & Barto, 2018). In this framework, one formulates the problem of interest as a Markov Decision Process (MDP), which can be solved in a variety of ways. In the RL paradigm, an agent interacts by means of actions with an unknown environment, receiving rewards that are proportional to the optimality of its actions; the objective of the agent is to adjust its behavior so as to maximize the sum of rewards received.
@@ 2.3.2 Dimensions of RL Algorithms
One important distinction is that between model-based algorithms (which assume access to a true or estimated model of the MDP) and model-free algorithms (which require only samples of agent-environment interactions). To be specific, the state space $\mathcal{S}$ and action space $\mathcal{A}$ are assumed to be known; a model $\mathcal{M}=(P,R)$ refers to knowing, or having some estimate of, the transition and reward functions $P,R$. Model-based methods can incorporate knowledge about the world to greatly speed up learning. The model can either be given a priori or learned. In the former case, it typically takes the form of a set of mathematical descriptions that fully define $P$ and $R$. In the latter case, learning $P$ corresponds to a density estimation problem, while learning $R$ is a Supervised Learning problem. Learning architectures for this purpose can range from probabilistic models such as Gaussian Processes (Deisenroth et al., 2011) to deep neural networks (Oh et al., 2015). Model-based RL is especially advantageous where real experience is expensive to generate and executing poor policies may have a negative impact (e.g., in robotics). When equipped with a model of the world, the agent can plan its policy, either at the time actions need to be taken focusing on the current state (decision-time planning), or not focused on any particular state (background planning). Model-free algorithms, on the other hand, can yield simpler learning architectures. This comes at the expense of higher sample complexity: they require more environment interactions to train. The two categories can also be combined: it is possible to use a model to generate transitions, to which a model-free algorithm can be applied (Gu et al., 2016). It is worth noting that, in the context of graph combinatorial optimization, we can often describe the transition and reward functions analytically based on the process of interest. For the TSP, both the transition and reward functions can be expressed in closed form and do not involve any uncertainty induced by the environment or other actors. Hence, in case the underlying model $\mathcal{M}$ is known, model-based algorithms can be leveraged.